If oil flows out of a pipe at the speed of 2.65m/s and the diameter of the pipe is 5.0cm, determine the discharge rate?

1
Expert's answer
2021-12-19T18:27:52-0500

radius of the pipe=5/2=2.5cm=0.025m5/2=2.5cm=0.025m

cross-section area of the pipe=πr2=3.142×0.0252=0.001964m2=\pi r^2=3.142\times 0.025^2 =0.001964m^2

discharge rate=speed of flow ×\times cross-section area=2.65×0.001964=0.005205m3/s2.65\times 0.001964=0.005205m^3/s